The coil consist of 90 ohm resistance and 127.3 mH inductance. Find the frequency of it when the impedance in the circuit is 50 ohm. 
(A) 75 Hz (B) 83 Hz (C) 50 Hz (D) 100Hz

Grade:12th pass

1 Answers

Vikas TU
5 years ago
This is a simple L-R circuit.
And the net impedence of such circuit is given b:
Z= sq.root(90^2 +  (w*127.3*10^-3)^2) = 50
sq both sides and solving we get,
2500 = 8100 + 0.162w^2
0.162w^2 = – 5600
the frequcny is coming negative which is not possible.
Hence the given options contadicts.